OpenStudy (anonymous):

Integral Calculus (Indefinite) what pointers should i think to choose the correct u. in u-substitution

zepdrix (zepdrix):

You're looking for a suitable `u` and `u'` to replace all of your integral.\[\Large \int\limits 3x^2\sqrt{x^3+4}\;dx\]Sometimes it helps to group things. This is how I would group them.\[\Large \int\limits\sqrt{x^3+4}\left(3x^2\;dx\right)\]

zepdrix (zepdrix):

In the example:\[\Large u=x^3+4 \qquad\qquad\qquad du=\left(3x^2\;dx\right)\]

zepdrix (zepdrix):

\[\Large \int\limits\limits\sqrt{x^3+4}\left(3x^2\;dx\right) \qquad\to \qquad \int\limits \sqrt{u}(du)\]

zepdrix (zepdrix):

It's really takes a bit of practice to recognize what is going on. Like you'll get a lot of weird looking stuff thrown at you. Sometimes it's hard to identify your du. Example:\[\Large \int\limits\frac{\ln x}{x}dx\]Another problem where I think grouping helps.\[\Large \int\limits\ln x\left(\frac{1}{x}dx\right)\]

zepdrix (zepdrix):

So in this example:\[\Large u=\ln x \qquad\qquad\qquad du=\left(\frac{1}{x}dx\right)\]

zepdrix (zepdrix):

It's not always that straight forward though. Sometimes you need to apply a `u-sub` and then mess around with the pieces. Example:\[\Large \int\limits \frac{x}{\sqrt{x-2}}dx\]

zepdrix (zepdrix):

In this example:\[\Large u=x-2 \qquad\qquad\qquad du=dx\]But we still need to deal with the x on top right? So we'll solve our initial substitution for x.\[\Large u=x-2 \qquad\to\qquad x=u+2\]

zepdrix (zepdrix):

\[\Large \int\limits\limits \frac{x}{\sqrt{x-2}}dx \qquad=\qquad \int\limits \frac{u+2}{\sqrt u}du\]This is much easier to solve in u now.

zepdrix (zepdrix):

Do you have any specific examples you need help with? :o Or any clarification on what I did above?

OpenStudy (dumbcow):

@zepdrix , very thorough examples :)

OpenStudy (anonymous):

cool now i get it its clear :), but what about the e^x and e^3x should i use the whole e^x as u? or just x?

zepdrix (zepdrix):

For a problem like this?\[\Large \int\limits e^{3x}dx\]Or something a bit more complex?

OpenStudy (anonymous):

something like this \[\int\limits \frac{ e ^{2x} }{ (1+e ^{x}) ^{1/3}} dx\]

OpenStudy (anonymous):

u see there are 2 \[e ^{x}\] but the other one is \[e ^{2x}\] how do i deal with this?

zepdrix (zepdrix):

mm sec XD thinking hehe

zepdrix (zepdrix):

Ok you have to think back to your rules of exponents. We're allowed to rewrite the numerator like this,\[\Large e^{2x}\quad=\quad (e^x)^2\]

zepdrix (zepdrix):

I'm gonna do my fun grouping thing again and maybe it will jump out at you. If not, that's ok. This one is a bit tricky.

OpenStudy (anonymous):

ok im at it, ill do my best

zepdrix (zepdrix):

\[\Large \int\limits\frac{e^x}{ (1+e ^{x}) ^{1/3}}\left(e^x\;dx\right)\]

zepdrix (zepdrix):

So I split up the square which allowed me to group it a little weird.

OpenStudy (anonymous):

haha never thought of doin that

OpenStudy (anonymous):

but what sould i use as a U

zepdrix (zepdrix):

Let's try letting the inside of our denominator be our u and see what happens. We might hit a bump in the road, if we do just have to think for a sec :) \(\Large u=1+e^x\)

zepdrix (zepdrix):

So what do you get for du? :x

OpenStudy (anonymous):

e^x

zepdrix (zepdrix):

\[\Large du=(e^x\;dx)\]Ok good. Hmm looks like we still need to deal with that e^x in the numerator. Any ideas? :) Look back at the last example I gave, with the sqrt in the denominator!! :O

OpenStudy (anonymous):

\[\int\limits \frac{(du)(du) }{(u)^{1/3} } dx\]

OpenStudy (anonymous):

oops XD

OpenStudy (anonymous):

this is tricky

zepdrix (zepdrix):

No no no :D we don't want multiple `differentials`. We just want 1 du! And remember that you're `replacing (e^x dx) with du. So that dx shouldn't be there anymore :O

zepdrix (zepdrix):

This is what we have so far,\[\Large \int\limits\limits\frac{e^x}{ (1+e ^{x}) ^{1/3}}\left(e^x\;dx\right) \qquad\to\qquad \int\limits\limits\frac{e^x}{ (u) ^{1/3}}\left(du\right)\]

zepdrix (zepdrix):

\[\Large \int\limits\limits\limits\frac{\color{#CC0033}{e^x}}{ (u) ^{1/3}}\left(du\right)\]

zepdrix (zepdrix):

Let's think back to our substitution that we set up,\[\Large u=1+\color{#CC0033}{e^x}\]

OpenStudy (anonymous):

i see wait let me continue the next step

OpenStudy (anonymous):

\[\int\limits \frac{u+1 }{ u ^{1/3} } du\]

OpenStudy (anonymous):

should be nagative i think

zepdrix (zepdrix):

oh on top?

OpenStudy (anonymous):

u-1 :D

zepdrix (zepdrix):

\[\Large \int\limits\limits\limits\limits\frac{u-1}{u^{1/3}}\left(du\right)\] Ok good good good :)

zepdrix (zepdrix):

From there we can split it into a couple of fractions. Remember how to do that? :o

zepdrix (zepdrix):

With a problem like this, after you make your u-sub and get everything plugged in, it might not `look` simpler, but it is! We have our addition/subtraction in the `numerator` now. So we can take advantage of fraction math :O

OpenStudy (anonymous):

yes \[\int\limits \frac{ u }{u ^{1/3}}-\frac{ 1 }{ u ^{1/3} } du\]

zepdrix (zepdrix):

cool, what next? :3

OpenStudy (anonymous):

\[\int\limits \frac{ u }{ u ^{1/3} } - \int\limits \frac{ 1 }{u ^{1/3} } du\]

zepdrix (zepdrix):

sure, if you wanna write it that way, that's fine

zepdrix (zepdrix):

divide your u's in the first int :O

OpenStudy (anonymous):

um theres an exponent on denominator u

OpenStudy (anonymous):

will subtraction of exponent work?

zepdrix (zepdrix):

ya that sounds right.

OpenStudy (anonymous):

\[\int\limits u ^{2/3} - \int\limits \frac{ 1 }{u ^{1/3} } du\]

OpenStudy (anonymous):

what do i do with \[\int\limits \frac{ 1 }{ u ^{1/3} }\]

zepdrix (zepdrix):

\[\Large \frac{1}{x^2}\quad=\quad x^{-2}\]

OpenStudy (anonymous):

haha then \[\int\limits u ^{-1/3}du do?\]

OpenStudy (anonymous):

How did you get do? O_o .

zepdrix (zepdrix):

do? :o

OpenStudy (anonymous):

wrote that @ wrong box :D

zepdrix (zepdrix):

oh haha

OpenStudy (anonymous):

Alright. It's right then :P .

OpenStudy (anonymous):

cool thanks now i just need to substitute u and make du a +c right?

zepdrix (zepdrix):

\[\Large \int\limits u^{2/3}-u^{-1/3}du\]Ya looks good so far :)

zepdrix (zepdrix):

Woops we still need to integrate !:O

OpenStudy (anonymous):

Well I woudn;t substitute yet.

OpenStudy (anonymous):

I would integrate it first and THEN substitute :P . NEVER substitute until the end haha.

OpenStudy (anonymous):

\[\int\limits_{?}^{?} \frac{ u ^{5/3} }{5/3 } - \int\limits_{?}^{?}\frac{ u ^{2/3} }{ 2/3 } du\]

OpenStudy (anonymous):

Well since you integrated the equation you can get rid of the intergal sign :P . There isn't any need of them since you solved the intergal.

OpenStudy (anonymous):

But yep that's fine :) .

OpenStudy (anonymous):

Now just simplify the thing :) .

OpenStudy (anonymous):

ok cool i got the answer tnx everyone :0

OpenStudy (anonymous):

Don't forget to add a +C!!

OpenStudy (anonymous):

btw how did u know again that u = 1+e^x and not the other one,

zepdrix (zepdrix):

Hmm that's a good question. Practice a little bit. If you're ever unsure, always try the denominator first D:

OpenStudy (anonymous):

Well you tend to pick u as something that it's derivative is also inside the intergal somewhere.

OpenStudy (anonymous):

But ultimately it comes down to practice.

OpenStudy (anonymous):

if i use u = e^x will it be ok? since i see there will be same du.

OpenStudy (anonymous):

How would you deal with the e^x +1 in the demonitor then?

zepdrix (zepdrix):

If you let u=e^x you would have ended up with something like this, \[\Large \int\limits \frac{u}{(1+u)^{1/3}}du\]

zepdrix (zepdrix):

Which unfortuately requires another substitution to solve easily.

OpenStudy (anonymous):

And personally fraction exponents are quite hard to integrate :3 .

OpenStudy (anonymous):

Often times in integrals you might have to use substitution more than once :P .

OpenStudy (anonymous):

like integration by parts?

OpenStudy (anonymous):

Maybe :P . It depends on the intergal.

OpenStudy (anonymous):

Intergation is basically an experiment.

OpenStudy (anonymous):

There may be multiple answers as well! :P .

OpenStudy (anonymous):

cool can i ask another topic of integral(trigo sub) here? or should i close this and post a new one?

OpenStudy (anonymous):

GO ahead :P .

zepdrix (zepdrix):

close it uppppp, it's getting long. Too much code, makes it laggy :(

OpenStudy (anonymous):

guess so :D